### 15 Apr 2026 v1

# A Case Study on Energy-Efficient Edge AI Crack Segmentation

Abstract
Crack segmentation on edge devices can support continuous infrastructure monitoring and maintenance and thereby help to preserve public safety. Furthermore, autonomous infrastructure monitoring by using Unmanned Aerial Vehicles (UAVs) can reduce inspection risks, as human operators no longer need to enter hazardous areas. Edge processing reduces the cost of inspection by eliminating the need for high resolution image storage for offline processing and mitigates the security risks and bandwidth requirements of streaming to cloud servers. Edge inference is difficult due to the limited memory and computational capabilities of edge devices, which can affect both accuracy and latency. Furthermore, battery-powered devices are subject to strict power and energy constraints. Together, these limitations impose restrictions on the model size and computational complexity that can be deployed close to the sensor. In recent years, Transformers have achieved state-of-the-art accuracy in a variety of applications, including semantic segmentation. However, Transformer-based models are typically large and computationally intensive, making efficient edge deployment difficult. To address this, we first apply knowledge distillation to enhance the performance of the base models. We then use Post-Training Quantization (PTQ) to compress the models further. Additionally, we consider the deployment of these models across multiple edge platforms. To maximize energy efficiency, we design and implement a custom hardware architecture for the models on an FPGA. Our results show that Knowledge Distillation (KD) improves all tested U-Net variants. Among the evaluated platforms, the selected FPGA implementation achieves $398$ Frames Per Second (FPS) at $204.99$ Frames/J while maintaining a mean Intersection over Union (IoU) of $69.42\%$. In addition, our best model reaches $71.92\%$ mean IoU, which is $8.82$ percentage points (pps) higher than the previously reported result on the CrackVision12K dataset.
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
I Introduction
Inspection and maintenance of civil infrastructure, such as bridges, tunnels, and high-rise buildings, are essential to ensure long-term safety and reliability. Traditional inspection methods are largely manual, requiring human experts to access potentially dangerous locations to visually assess structural damage. With the emergence of UAVs, automated large-scale inspection of concrete and asphalt surfaces has become feasible, offering both safety and efficiency advantages. UAV-based inspection systems equipped with onboard computer vision can identify cracks and other forms of surface degradation without exposing human inspectors to hazardous environments.
The main challenge for onboard crack segmentation is the need to operate on resource-constrained edge devices with limited power, memory, and computational capacity, while still enabling real-time processing. Although Transformer-based models achieve state-of-the-art segmentation performance, their computational and memory demands hinder practical edge deployment. Consequently, lightweight architectures such as U-Net remain more suitable. Despite their simpler design, U-Net variants exhibit substantial variability in parameter count and computational cost, and still require optimization for efficient edge execution. Identifying models that balance segmentation accuracy and efficiency is therefore critical for real-world edge crack segmentation.
Efficient deployment of crack segmentation models on edge devices is governed by strict constraints on latency, power consumption, and energy efficiency, particularly under small batch sizes dictated by the number of input sensors (e.g., cameras). In this work, we investigate implementation across heterogeneous edge platforms, including CPUs, GPUs, and Neural Processing Units (NPUs), each offering distinct trade-offs. GPUs provide high throughput, especially for large batch sizes, but typically incur higher power consumption and are less efficient in small-batch regimes. CPUs offer greater flexibility and broader operator support, yet often fall short in terms of parallel efficiency. NPUs, while promising superior energy efficiency, are constrained by limited support for operators and model architectures, complicating deployment. These disparities make platform selection, numerical precision, and runtime configuration non-trivial design choices. Collectively, these limitations highlight the need for custom Field-Programmable Gate Array (FPGA)-based architectures to achieve superior energy efficiency and mitigate platform-specific constraints.
Our contributions are as follows:
  • We perform a comprehensive design space exploration across multiple levels, from model-level design to hardware platforms, providing a comparative analysis of segmentation results, latency, power consumption, and energy efficiency for crack segmentation on several edge platforms.
  • We design and implement U-Net variants spanning three orders of magnitude in parameter count to study the trade-off between segmentation performance and efficiency. Knowledge distillation from a Transformer model is applied to enhance the smaller U-Net variants. The resulting Pareto-optimal model is $63.63\times$ smaller, and 6.32pps better, compared to the currently best tested Hybrid-Segmentor model from Goo et al [11].
  • To further improve energy efficiency and throughput for single-batch processing, we develop a custom hardware architecture and implement it on an FPGA.
  • The Pareto-optimal FPGA implementation outperforms alternative embedded platforms, achieving the highest energy efficiency (204.99 Frames/J) and throughput (398 FPS) while still improving the segmentation accuracy by 2.79pps compared to the baseline U-Net without KD.
II Related Work
II-A Crack Segmentation Approaches
Early methods of crack segmentation are often based on graph-based approaches such as the use of shortest path algorithms, as in [12], or the use of minimum spanning trees (MST), as in [35]. The advantages of these algorithms lie in their efficient implementation and explainability. However, since the introduction of UNet [25], deep learning models recognize cracks much better. Due to its efficient encoder-decoder structure, skip connections, UNet has become particularly popular for crack segmentation, especially for real-time applications. As a result, there are now many deep learning approaches that build on this and propose new, improved models or training procedures [6, 20, 14, 19, 29, 18]. UNet++ was introduced in 2019 [34]. In 2023, an improved UNet incorporating an attention mechanism was proposed [1]. However, most of these variants require increased computational resources and are therefore unsuitable for edge or embedded platforms. One of the most advanced real-time capable models is UNext [2]. However, its reliance on attention mechanisms introduces architectural complexity that currently hinders edge deployments.
II-B Semantic Segmentation on Embedded Platforms
In recent years, embedded platforms have gained significant traction for crack segmentation. Some authors propose specifically designed segmentation models for embedded platforms. For example, [3] presents MCFF-L Net - an architecture with 1.18 million parameters - combined with knowledge distillation, which achieves good segmentation results and is able to run in real-time on a Jetson Xavier NX. Similarly, [33] considered crack segmentation on devices for structural health monitoring. A related approach by [32] uses a special KD approach for measurement tasks on embedded systems.
Other works focus on simplifying UNet for microcontroller-class hardware, like [8], which analyzed small CNN-based crack segmentation models on low-power microcontrollers. In contrast, [9] showed that UNet models can be compressed to meet the requirements of very resource-constrained devices such as the STM32 platform.
II-C Semantic Segmentation on FPGA
Recently, FPGAs have become a popular choice for implementing semantic segmentation algorithms under real-time and limited power budget constraints. In [10], a compressed ENet is deployed entirely on-chip using HLS4ML, while [17] and [26] combine pruning, quantization, and Vitis-AI flow to run encoder-decoder networks in real time on FPGA. Other approaches redesign segmentation blocks specifically for hardware efficiency, such as [4]. Another approach to shrinking the model is about model compression and low-bit quantization. In [23], DeepLabV3+ is pruned using a hardware-aware genetic algorithm, while in [22] a UNet is quantized to 3-bit precision and implemented on a dedicated on-chip pipeline. Similar OpenCL- or HLS-based encoder-decoder accelerators further show the impact of quantization and optimized data flow on performance [31]. However, in the context of crack segmentation, FPGA-based solutions are less common. The work in [5] represents one of the few examples, implementing a particle-filter–based crack detector directly on an FPGA for real-time inspection. Although they use a totally different approach than we do, their approach already shows the relevance of FPGAs for building inspections.
II-D FPGA-based U-Net Implementations
Existing work on U-Net deployment on FPGAs can be broadly categorized into two architectural paradigms: reconfigurable (engine-based) architectures and fully streaming (dataflow) architectures.
Reconfigurable architectures instantiate a limited set of generic compute engines in the Programmable Logic (PL), which are time-multiplexed across network layers. Each layer is executed by reprogramming the compute engines with the corresponding parameters and supplying the required input feature maps. This approach, explored in [24], [16], and [13], provides high flexibility and enables the deployment of networks whose size exceeds on-chip resource capacity. However, it incurs frequent off-chip memory transfers for intermediate feature maps and parameters, leading to increased memory bandwidth requirements and higher energy consumption.
In contrast, streaming architectures map the entire network onto the PL as a sequence of dedicated processing modules, enabling pipelined execution with minimal intermediate storage. This dataflow-oriented design significantly reduces memory traffic and improves energy efficiency by exploiting on-chip data reuse. However, the approach is constrained by the available on-chip resources, particularly block RAM and logic, which must accommodate both model parameters and intermediate activations.
U-Net architectures pose additional challenges for streaming implementations due to their encoder–decoder structure with multiple skip connections that require buffering high-resolution feature maps. These skip connections substantially increase on-chip memory demand and can become the primary bottleneck in fully unrolled designs based on widely used FINN [28] and HLS4ML [7] hardware libraries. To address this limitation, we propose offloading skip connection feature maps to external memory. While this introduces additional memory accesses, the resulting access pattern is largely sequential and predictable, allowing efficient utilization of memory bandwidth. Consequently, the associated energy overhead is expected to be limited, while significantly relaxing on-chip memory constraints and enabling the deployment of larger U-Net variants.
III Methodology
III-A Overview
Our objective is to enable efficient crack segmentation on resource-constrained edge devices, including embedded AI accelerators and FPGAs, without compromising accuracy. We systematically derive a family of U-Net variants by scaling the number of feature channels, creating models spanning three orders of magnitude in parameter count. To mitigate the performance degradation of smaller models, we leverage knowledge distillation from a high-capacity Transformer teacher, allowing compact U-Nets to retain segmentation accuracy close to the baseline. This approach, combined with hardware-aware optimization, facilitates the identification of Pareto-optimal models that balance efficiency and performance across diverse edge platforms.
III-B U-Net Channel Scaling
We use the original UNet model [25] as a basis and generate four smaller variants by reducing the number of feature maps. For this purpose, we introduce a parameter $c\in\{2,4,8,16,32\}$ which controls the number of all feature maps in the network while preserving the original encoder–decoder architecture. Setting $c=32$ reproduces the original UNet architecture. This parametrization allows us to analyze how the number of feature maps influences the segmentation results as well as the runtime and power consumption.
Each encoder stage consists of two $3\times 3$ convolutions with padding $1$, each followed by Batch Normalization and a ReLU activation. Downsampling is done by using a $2\times 2$ max-pooling. The decoder mirrors this structure, using either nearest upsampling (only for Edge TPU) or learnable $2\times 2$ transposed convolutions. After upsampling, the decoder concatenates the feature maps with the corresponding encoder activations via skip connections and also applies two $3\times 3$ convolutions with padding $1$, where each is followed by Batch Normalization and a ReLU activation. Finally, the network uses a $1\times 1$ convolution to map the feature representation to the two target classes (crack and background). All convolutions in the model are implemented without bias terms, as each is directly followed by a Batch Normalization layer.
III-C Dataset and Preprocessing
In this work, we use the CrackVision12K dataset introduced by Goo et al. [11]. CrackVision12K is a curated benchmark for crack segmentation that combines images from several previously published crack datasets into a unified dataset with improved and standardized annotations. Compared with the original source datasets, the labels were manually refined to reduce annotation noise and inconsistencies. In addition, all images and masks were resized to a common spatial resolution of $256\times 256$ pixels.
In total, the dataset contains 12000 image-mask pairs. To ensure comparability with prior work on CrackVision12K, we use the train/validation/test split provided by Goo et al. [11]. The training set contains 9600 images, while the validation and test sets each contain 1200 images.
III-D Data Augmentation
We also applied data augmentation methods during training to further improve the segmentation results and make the models more robust for practical applications. Each augmentation is applied with a probability of 50%. The augmentation pipeline includes additive image noise, motion blur, horizontal flipping, and small random rotations.
III-E Knowledge Distillation
To improve the segmentation results even more, we use KD, where a stronger teacher network guides the training of a smaller student network [15]. In our setup, the teacher is a SegFormer-B5 model [30] that is first fine-tuned on CrackVision12K. The student is one of our channel-scaled U-Net variants. During KD training, the teacher remains frozen, while the student is optimized.
Let $x\in\mathbb{R}^{H\times W\times 3}$ denote an input image and $y$ its ground-truth segmentation mask. The teacher produces logits $z_{t}(x)$ and the student produces logits $z_{s}(x)$. Since the spatial resolution of the teacher and student models is not equal, we resize the teacher logits to match the student’s output resolution by using bilinear interpolation. Afterwards, we apply KD on the softened class distributions:
$p_{t}(x)=\mathrm{softmax}\left(\tfrac{z_{t}(x)}{T}\right),\hskip 5.69054ptp_{s}(x)=\mathrm{softmax}\left(\tfrac{z_{s}(x)}{T}\right)$ (1)
where $T>1$ is the temperature parameter.
The KD loss is defined as the Kullback-Leibler divergence between teacher and student predictions:
$\mathcal{L}_{KD}=T^{2}\cdot\mathrm{KL}\!\left(p_{t}(x)\,\|\,p_{s}(x)\right).$ (2)
In addition to the soft teacher supervision, we use a hard loss based on the ground-truth labels. More specifically, we combine weighted cross-entropy with a soft Dice loss computed on the crack class:
$\mathcal{L}_{hard}=\mathcal{L}_{CE}+\lambda_{Dice}\,\mathcal{L}_{Dice}$ (3)
The factor $\lambda_{Dice}$ weights the impact of the Dice loss and is determined by our hyperparameter search. Thus, our final training loss is then:
$\mathcal{L}=\alpha\mathcal{L}_{KD}+(1-\alpha)\mathcal{L}_{hard},$ (4)
where $\alpha\in[0,1]$ controls the trade-off between teacher supervision and supervision from the ground-truth labels.
III-F Training Setup and Hyperparameter Search
We first train all five U-Net variants by using supervised-learning to get the baseline models. In this setup, we use a combination of random search and an evolutionary algorithm as a hyperparameter search, adopted from Tschöpe et al. [27]. The optimized hyperparameters include the initial learning rate of the AdamW optimizer [21], the learning rate scheduler decay factor $\gamma$, the scheduler step size, the momentum values of the AdamW optimizer, the weight decay, and the class-weights that are used in the cross-entropy loss. However, for testing we use fixed class-weights, computed by the pixel distribution. We explain this in more detail in section III-G.
For training with KD, we additionally optimize the KD temperature $T$, the KD mixing factor $\alpha$, and the weighting factor of the Dice term $\lambda_{Dice}$. The hard supervision consists of weighted cross-entropy and a soft Dice loss on the crack class, while the soft supervision is given by the KL divergence between teacher and student predictions.
All models are trained for at most 150 epochs with early stopping. For each model configuration, we repeat the training with five random seeds and report the mean and standard deviations of the evaluation metrics.
III-G Evaluation Metrics
We report both mean IoU and a weighted IoU. Let $\mathrm{IoU}_{bg}$ and $\mathrm{IoU}_{crack}$ denote the IoU values for the background and crack classes, respectively. The mean IoU is defined as
$\mathrm{mIoU}=0.5\cdot\mathrm{IoU}_{bg}+0.5\cdot\mathrm{IoU}_{crack}$ (5)
Thus, this metric weights the background pixels and crack pixels equally by a factor of $0.5$. Since crack pixels are much rarer but more important than background pixels in our application, we additionally report a weighted IoU:
$\mathrm{wIoU}=w_{bg}\cdot\mathrm{IoU}_{bg}+w_{crack}\cdot\mathrm{IoU}_{crack},$ (6)
with
$w_{bg}=0.068,\qquad w_{crack}=0.932.$ (7)
The weights are computed from the inverse class frequencies of the training set and normalized to sum to one. As a result, the weighted IoU emphasizes the crack class much more strongly than the background class. Therefore, a model can achieve a high mean IoU due to very good background segmentation while still obtaining a noticeably lower weighted IoU if its crack segmentation remains challenging.
IV Implementation
This work investigates the performance of efficient U-Net implementations on a diverse range of embedded computing platforms. Besides FPGA, the comparison encompasses conventional CPU, GPU, and a dedicated AI accelerator, allowing for a comprehensive evaluation of deployment strategies under different computational and energy constraints.
IV-A Edge Platforms
Specifically, we target: the Raspberry Pi 5 Model B (8 GB) featuring a quad-core Cortex‑A76 CPU, the Seeed Studio reComputer J3011 equipped with an NVIDIA Jetson Orin Nano (8 GB), the Coral Edge TPU M.2 module hosted on a Raspberry Pi 5. For CPU-based execution, the U-Net models were run using the ONNX Runtime with floating-point weights and activations. On the GPU platform, two inference configurations were explored: one fully relying on floating-point precision and the other employing int8 quantization for network parameters while retaining floating-point activations. The GPU experiments were conducted using the TensorRT execution engine to maximize efficiency. On the dedicated AI accelerator, the TensorFlow Lite models, quantized to int8 for both weights and activations, were compiled using the Edge TPU compiler. Inference latency was measured for batch size 1 and reported as a mean value of five execution runs on a subset of 1024 input images. The reported latency includes data transfer times between the host and the device, as well as the time required for conversions between fp32 and int8 data formats for inputs and outputs, respectively. Power consumption was monitored throughout the experiments using an external USB POWER‑Z KM003C KT002 multimeter. Both idle and runtime conditions were evaluated, with the latter involving continuous inference execution for a duration of 20 seconds. The mean power values were derived from three independent measurements. For the Coral Edge TPU setup, power readings correspond to the complete configuration, capturing the consumption of both the M.2 module and the Raspberry Pi 5 host. Active cooling solutions were employed for all Raspberry Pi devices, while the Coral Edge TPU operated with an aluminum heat sink to ensure thermal stability during execution.
IV-B Custom Hardware Implementation on FPGA
The FPGA implementation is built upon a High-level Synthesis (HLS) hardware library that provides custom hardware architectures for a wide range of Deep Neural Network (DNN) layers and components. These architectures are highly customizable, enabling the implementation of diverse neural topologies and facilitating rapid design-space exploration. The hardware library is implemented as a collection of C++ template functions equipped with HLS annotations. The overall architecture is optimized for low-power consumption and ultra-low latency operation. This is primarily accomplished by storing all weights and the majority of intermediate results in on-chip memory, as off-chip data transfers incur higher energy costs and introduce additional latency. Nevertheless, a design relying solely on on-chip resources inherently limits the size of DNN models that can be feasibly implemented. To overcome this constraint, we developed a hybrid hardware architecture that supports off-chip implementation of skip and residual connections, effectively relaxing the limitations on model size and structural complexity. The architecture is fully pipelined, allowing all layers to operate concurrently and initiate computation immediately once their inputs become available, thereby reducing both latency and energy consumption. Individual hardware modules corresponding to each layer are interconnected via on-chip data streams and assembled into a single top-level module, as illustrated in Fig. 1.
Figure 1: The sample hardware architecture where each layer is mapped to a separate hardware instance. The modules are connected via on-chip data streams.
Power was measured using the onboard Texas Instruments IN219 current/power monitor. Idle power was measured after configuring the PL for a continuous period of 10 seconds. While the runtime power was averaged over 32 repetitions, each consisting of 32 images processed in series. The measurement includes the PL reading the image from the memory, processing it, and writing the result back to the memory. The same methodology was used for measuring latency.
V Results
V-A Accuracy–Efficiency Trade-off Analysis
Table I summarizes the segmentation performance of all models tested in the CrackVision12K dataset before quantization. All models shown in Table I were trained on five different random seeds. Therefore, the mean and standard deviation for the weighted IoU and the unweighted IoU are provided.
Considering the model results without knowledge distillation, the original U-Net ($c=32$) serves as the baseline. Reducing the base channel count to $c=16$ reduces the number of Multiply-Accumulate (MAC) operations by approximately $4\times$, from 24.21G to 6.10G, while the accuracy drop is negligible - only 0.18pps in weighted IoU and 0.08pps in mean IoU. Reducing further to $c=8$ provides an additional $\sim 4\times$ MAC reduction to approximately 1.55G, with a still modest accuracy cost of only 3.63pps in weighted IoU and 2.11pps in mean IoU relative to the baseline - a competitive result given the overall $16\times$ reduction in computational cost. Scaling down to $c=2$, however, incurs steep accuracy costs: a drop of 13.60pps in weighted IoU and 10.46pps in mean IoU compared to $c=16$, confirming that $c=16$ and $c=8$ represent the best efficiency–accuracy trade-offs in this family. As expected, SegFormer B5 outperforms all UNet variants without knowledge distillation, owing to its considerably larger capacity and modern transformer-based architecture.
With knowledge distillation, all UNet variants improve. The baseline model ($c=32$) achieves a weighted IoU of $51.06\%$ and a mean IoU of $71.92\%$—within 2.07pps of the teacher SegFormer B5 ($53.13\%$ weighted IoU), showing that KD brings the student nearly to the teacher’s level. Crucially, the efficiency gains identified without KD are preserved: $c=16$ achieves $50.63\%$ weighted IoU and $71.74\%$ mean IoU, only 0.43 and 0.18pps below $c=32$, while still benefiting from the same $4\times$ MAC reduction. Similarly, $c=8$ achieves $49.25\%$ weighted IoU and $70.96\%$ mean IoU—only 1.81pps and 0.96pps below the best UNet model —at a $16\times$ overall reduction in compute.
Overall, these results show that KD not only improves segmentation results across all UNet variants, but also makes the models significantly more robust to channel reduction: the mean IoU gap between $c=32$ and $c=8$ shrinks from 2.19pps without KD to only 0.96pps with KD, which shows that KD compensates for the capacity loss introduced by pruning. In the next subsection, we evaluate each variant across several edge platforms under quantization.
TABLE I: Comparison of U-Net variants with different channel scaling factors, with and without knowledge distillation (KD). The Hybrid-Segmentor model was proposed by Goo et al [11]. SegFormer B5 is included as a high-capacity reference.
Approach
Weighted
IoU[%]
Mean
IoU[%]
GOPs
Num.
Params (M)
Hybrid-Segmentor - $63.1\pm 0.36$ - -
SegFormer B5 $53.13\pm 0.21$ $73.42\pm 0.26$ $51.25$ $84.59$
Base 2 w/o KD $29.31\pm 0.54$ $56.17\pm 0.43$ $0.11$ $0.031$
Base 4 w/o KD $37.67\pm 0.43$ $62.54\pm 0.32$ $0.40$ $0.122$
Base 8 w/o KD $39.10\pm 1.13$ $64.44\pm 0.81$ $1.55$ $0.487$
Base 16 w/o KD $42.73\pm 1.50$ $66.55\pm 1.06$ $6.10$ $1.943$
Base 32 w/o KD * $42.91\pm 0.56$ $66.63\pm 0.43$ $24.21$ $7.763$
Base 2 KD $47.54\pm 0.31$ $69.90\pm 0.27$ $0.11$ $0.031$
Base 4 KD $48.81\pm 0.95$ $70.68\pm 0.78$ $0.40$ $0.122$
Base 8 KD $49.25\pm 0.67$ $70.96\pm 0.59$ $1.55$ $0.487$
Base 16 KD $50.63\pm 0.55$ $71.74\pm 0.44$ $6.10$ $1.943$
Base 32 KD $\mathbf{51.06\pm 0.71}$ $\mathbf{71.92\pm 0.65}$ $\mathbf{24.21}$ $\mathbf{7.763}$
  • * This model is used as a baseline.
V-B Hardware Implementation Results
Table IV shows the results for knowledge-distilled models across three embedded platforms and the FPGA, reporting segmentation accuracy (weighted and mean IoU), throughput (FPS), and power and energy metrics. In addition to idle and runtime power, both dynamic and runtime energy efficiency (frames/J) are provided. Our analysis focuses primarily on dynamic energy efficiency, as it isolates the energy consumption attributable to active inference. In contrast, runtime energy efficiency is affected by background processes and system-level overhead, making cross-platform comparisons less reliable. Fig. 2 further illustrates the trade-off between dynamic energy efficiency and mean IoU across the evaluated configurations.
V-B1 FPGA results
For an efficient FPGA implementation PTQ is used to reduce the precision of the model parameters. Two experiments were conducted; in the first experiment, the weights and activations were quantized to int8 as shown in Table II this experiment achieved a similar or higher IoU compared to the fp32 experiment due to regularization. In the second experiment, the weights were quantized to int4, which resulted in a considerable drop in accuracy as shown in Table II. We attribute this drop to PTQ and expect quantization aware training (QAT) to recover this drop. Based on the results in Table II int8 models were chosen for implementation on FPGA, for which the AMD Vitis 2024.2 HLS tool was used and the target platform was set to TySOM-2A-ZU19EG. Table III presents the resources required to implement the chosen models on the target platform. Two experiments were conducted; in the first experiment, the resource-heavy skip connections of the UNet model were kept on chip, while in the second experiment, they were moved to off-chip memory. This made it possible to implement the larger model with $c=16$. Skip connections were moved to off-chip memory by creating an AXI interface for each skip connection, which is used for reading and writing. This is required due to the high read and write speeds and the fact that all skip connections are accessed in parallel. On-chip buffers were added to the design to bridge the gap between the high-speed PL and the slow off-chip memory. The extra buffers and interfaces lead to an average increase of $2.19\%$ and $1.60\%$ in LookUp Tables (LUTs) and Flip-Flops (FFs) respectively, which is within the acceptable overhead. This shows the tradeoff between moving skip connections off chip and keeping them on chip, which is less efficient for small models i.e., with $c=2$. Table IV shows the throughput, IoU, energy and efficiency of the implemented models on the FPGA. All FPGA implementations achieved the same IoU as the corresponding software implementations.
V-B2 Edge Computing Platforms
[…]
Although the Coral Edge TPU supports transposed convolutions, models that rely on transposed convolution for upsampling experience a significant drop in accuracy when deployed on the Coral Edge TPU. To address this, the models were retrained using nearest upsampling, which preserved their accuracy. However, the model with base 32 failed to run on the Coral Edge TPU.
TABLE II: Post-training quantization results for models with knowledge distillation.
Weight
Precision
Base
Activation
Precision
Weighted
IoU[%]
Mean
IoU[%]
int8 2 8 45.49 68.91
4 8 46.29 69.42
8 8 47.86 70.35
16 8 49.22 71.11
32 8 49.05 70.92
int4 2 8 40.62 65.09
4 8 42.37 66.76
8 8 42.87 67.32
16 8 38.23 64.85
32 8 48.46 70.73
TABLE III: Implementation Resource Utilization (%) on AMD Zynq UltraScale+ ZU19EG-FFVB1517 FPGA.
Skip Base
LUTs
$F$s
BRAMs 36K
DSPs
On-chip 2 11.65 3.35 15.24 11.18
4 16.81 4.51 33.23 43.14
8 23.45 5.88 78.86 92.17
Off-chip 2 13.90 4.94 12.40 11.18
4 19.13 6.14 28.00 43.14
8 25.48 7.47 55.79 92.17
16 36.88 8.07 100.00 97.00
TABLE IV: Comparison of throughput ($\theta$), IoU, and energy efficiency for various model precisions on RPi5, Orin Nano, Edge TPU, and FPGA with skip connections mapped to on-chip and off-chip memory at batch size 1.
Base Device
Model
[bits]
Data
[bits]
Weighted
IoU
Mean
IoU
$\theta$
[FPS]
Idle
Power
[W]
Runtime
Power
[W]
Dynamic
Energy
Efficiency
[Frames/J]
Runtime
Energy
Efficiency
[Frames/J]
2 RPi5 fp32 fp32 45.28 68.78 246 2.73 9.74 35.09 25.26
Orin Nano fp32 fp32 45.28 68.78 378 6.97 9.98 125.58 37.88
int8 fp32 45.18 68.75 383 6.97 10.00 126.40 38.30
Edge TPU int8 int8 46.31 69.44 222 3.38 5.44 107.77 40.81
FPGA (on-chip) int8 int8 45.49 68.91 397 18.71 19.47 522.98 20.41
FPGA (off-chip) int8 int8 45.49 68.91 398 18.74 19.47 544.65 20.42
4 RPi5 fp32 fp32 47.32 69.86 98 2.73 9.69 14.08 10.11
Orin Nano fp32 fp32 47.32 69.86 329 6.97 10.77 86.58 30.55
int8 fp32 47.34 69.87 361 6.97 10.11 114.97 35.71
Edge TPU int8 int8 47.95 70.36 211 3.27 5.46 96.35 38.64
FPGA (on-chip) int8 int8 46.29 69.42 398 18.89 20.83 204.99 19.09
FPGA (off-chip) int8 int8 46.29 69.42 398 19.08 21.12 194.90 18.83
8 RPi5 fp32 fp32 47.56 70.13 34 2.73 9.93 4.72 3.42
Orin Nano fp32 fp32 47.56 70.13 247 6.97 11.56 53.81 21.37
int8 fp32 47.56 70.13 320 6.97 10.42 92.75 30.71
Edge TPU int8 int8 48.59 70.76 160 3.27 4.83 102.56 33.13
FPGA (on-chip) int8 int8 47.86 70.35 397 19.29 26.76 53.16 14.84
FPGA (off-chip) int8 int8 47.86 70.35 398 19.32 26.79 53.23 14.84
16 RPi5 fp32 fp32 48.50 70.59 12 2.73 10.54 1.54 1.14
Orin Nano fp32 fp32 48.50 70.59 158 6.97 12.60 28.06 12.54
int8 fp32 48.50 70.58 276 6.97 11.18 65.56 24.69
Edge TPU int8 int8 49.26 71.16 69 3.27 4.71 47.92 14.65
FPGA (off-chip) int8 int8 49.22 71.11 82 19.75 27.22 10.99 3.02
32 RPi5 fp32 fp32 49.22 70.98 4 2.73 11.06 0.48 0.36
Orin Nano fp32 fp32 49.22 70.98 78 6.97 13.88 11.29 5.62
int8 fp32 49.06 70.88 222 6.97 12.04 43.79 18.44
$10^{0}$$10^{1}$$10^{2}$$10^{3}$696969.569.5707070.570.57171Dynamic Energy Efficiency [frames/J]Mean IoU [%]Pareto frontBase 2Base 4Base 8Base 16Base 32RPi5Orin Nano fp32Orin Nano int8Edge TPUFPGA on-chipFPGA off-chip
Figure 2: Pareto frontier showing the trade-off between accuracy (weighted IoU) and energy efficiency across different platforms and model sizes. FPGA on-chip: models UNet skip connections stored on chip, FPGA off-chip: models UNet skip connections stored off chip in the memory
VI Conclusion and Future Work
In this work, we studied crack segmentation on resource-constrained edge platforms from two sides: The model design and the target hardware. For this, we scaled the U-Net architecture over three orders of magnitude in parameter count and analyzed the trade-off between IoU and computational cost. Our results show that reducing the model size can strongly lower the computational cost, while some of the smaller variants still remain close to the baseline in terms of segmentation accuracy. To improve our scaled UNet models further, we applied knowledge distillation by using SegFormer B5 as the teacher model. This improved all U-Net variants.
We then evaluated all variants on several edge platforms and compared their throughput, power consumption, and energy efficiency. We also developed a custom FPGA-based architecture for single-batch inference. Among all tested platforms, the FPGA (on-chip) implementation with the UNet Base 4 using KD and int8 quantization achieves an energy efficiency of 204.99 Frames/J and a throughput of 398 FPS, while improving the mean IoU by 2.79pps compared to the baseline model (UNet Base 32 without KD). Compared to the Hybrid-Segmentor model from Goo et al. [11], which is - as far as we know - currently the best tested model on this dataset, our aforementioned UNet Base 4 model with KD is 6.32pps better in mean IoU.
Therefore, our results show that the combination of model scaling, knowledge distillation, quantization, and hardware-specific implementation is useful for crack segmentation on edge devices.
For future work, we plan to investigate quantization-aware training and structured pruning to reduce the model cost further. In addition, it would be interesting to study other teacher models or multi-teacher settings for knowledge distillation. Finally, we plan to record and annotate our own crack segmentation dataset from UAV flights to evaluate the models under more realistic conditions.

### 06 May 2026 v2

@@
 Together, these limitations impose restrictions on the model size and computational complexity that can be deployed close to the sensor.
-In recent years, Transformers have achieved state-of-the-art accuracy in a variety of applications, including semantic segmentation.
-However, Transformer-based models are typically large and computationally intensive, making efficient edge deployment difficult.
-To address this, we first apply knowledge distillation to enhance the performance of the base models.
+Recent Transformer-based models reach high accuracy on semantic segmentation, but they are typically large and computationally intensive, which makes them hard to deploy on edge devices.
+We therefore use knowledge distillation to improve the segmentation results of smaller base models.
 We then use Post-Training Quantization (PTQ) to compress the models further.
@@
 In addition, our best model reaches $71.92\%$ mean IoU, which is $8.82$ percentage points (pps) higher than the previously reported result on the CrackVision12K dataset.
-This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
+©2026 IEEE.
+Personal use of this material is permitted.
+Permission from IEEE must be obtained for all other uses, including reprinting/republishing this material for advertising or promotional purposes, collecting new collected works for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
 I Introduction
 Inspection and maintenance of civil infrastructure, such as bridges, tunnels, and high-rise buildings, are essential to ensure long-term safety and reliability.
-Traditional inspection methods are largely manual, requiring human experts to access potentially dangerous locations to visually assess structural damage.
+Traditional inspection is mostly manual: human experts have to access dangerous locations to look for structural damage.
 With the emergence of UAVs, automated large-scale inspection of concrete and asphalt surfaces has become feasible, offering both safety and efficiency advantages.
@@
 The main challenge for onboard crack segmentation is the need to operate on resource-constrained edge devices with limited power, memory, and computational capacity, while still enabling real-time processing.
-Although Transformer-based models achieve state-of-the-art segmentation performance, their computational and memory demands hinder practical edge deployment.
-Consequently, lightweight architectures such as U-Net remain more suitable.
-Despite their simpler design, U-Net variants exhibit substantial variability in parameter count and computational cost, and still require optimization for efficient edge execution.
+Transformer-based models reach high segmentation accuracy, but their compute and memory demands make them difficult to run on edge devices.
+Lightweight architectures like U-Net are therefore the more practical choice.
+However, U-Net variants differ a lot in parameter count and compute cost, and still need to be optimized for edge use.
 Identifying models that balance segmentation accuracy and efficiency is therefore critical for real-world edge crack segmentation.
 Efficient deployment of crack segmentation models on edge devices is governed by strict constraints on latency, power consumption, and energy efficiency, particularly under small batch sizes dictated by the number of input sensors (e.g., cameras).
-In this work, we investigate implementation across heterogeneous edge platforms, including CPUs, GPUs, and Neural Processing Units (NPUs), each offering distinct trade-offs.
+In this work, we look at the implementation on different edge platforms, including CPUs, GPUs, and Neural Processing Units (NPUs), which all have different tradeoffs.
 GPUs provide high throughput, especially for large batch sizes, but typically incur higher power consumption and are less efficient in small-batch regimes.
@@
 Our contributions are as follows:
-  • We perform a comprehensive design space exploration across multiple levels, from model-level design to hardware platforms, providing a comparative analysis of segmentation results, latency, power consumption, and energy efficiency for crack segmentation on several edge platforms.
+  • We explore the design space over model scaling, numerical precision, and hardware platform, and we compare segmentation accuracy, latency, power, and energy efficiency for crack segmentation on four edge platforms (CPU, GPU, TPU, and FPGA).
   • We design and implement U-Net variants spanning three orders of magnitude in parameter count to study the trade-off between segmentation performance and efficiency.
-Knowledge distillation from a Transformer model is applied to enhance the smaller U-Net variants.
-The resulting Pareto-optimal model is $63.63\times$ smaller, and 6.32pps better, compared to the currently best tested Hybrid-Segmentor model from Goo et al [11].
+We use KD from a Transformer model (teacher) to improve the smaller U-Net model (student).
+Under the train/val/test split released in [12] for CrackVision12K, our most accurate model reaches 8.82pps higher mean IoU than their Hybrid-Segmentor model, while using $30.64\times$ fewer parameters.
   • To further improve energy efficiency and throughput for single-batch processing, we develop a custom hardware architecture and implement it on an FPGA.
-  • The Pareto-optimal FPGA implementation outperforms alternative embedded platforms, achieving the highest energy efficiency (204.99 Frames/J) and throughput (398 FPS) while still improving the segmentation accuracy by 2.79pps compared to the baseline U-Net without KD.
+  • Our Pareto-optimal FPGA implementation reaches the highest energy efficiency (204.99 Frames/J) and throughput (398 FPS) of all tested platforms, while also improving the mean IoU by 2.79pps over the baseline U-Net without KD.
 II Related Work
 II-A Crack Segmentation Approaches
-Early methods of crack segmentation are often based on graph-based approaches such as the use of shortest path algorithms, as in [12], or the use of minimum spanning trees (MST), as in [35].
-The advantages of these algorithms lie in their efficient implementation and explainability.
-However, since the introduction of UNet [25], deep learning models recognize cracks much better.
-Due to its efficient encoder-decoder structure, skip connections, UNet has become particularly popular for crack segmentation, especially for real-time applications.
-As a result, there are now many deep learning approaches that build on this and propose new, improved models or training procedures [6, 20, 14, 19, 29, 18].
-UNet++ was introduced in 2019 [34].
-In 2023, an improved UNet incorporating an attention mechanism was proposed [1].
-However, most of these variants require increased computational resources and are therefore unsuitable for edge or embedded platforms.
-One of the most advanced real-time capable models is UNext [2].
-However, its reliance on attention mechanisms introduces architectural complexity that currently hinders edge deployments.
-II-B Semantic Segmentation on Embedded Platforms
-In recent years, embedded platforms have gained significant traction for crack segmentation.
-Some authors propose specifically designed segmentation models for embedded platforms.
+Early methods of crack segmentation are often based on graph-based approaches, such as shortest path algorithms [13] or minimum spanning trees [39].
+These methods are efficient and explainable, but since the introduction of U-Net [28], deep learning models recognize cracks much better.
+Due to its efficient encoder-decoder structure with skip connections, U-Net has become particularly popular for crack segmentation, especially for real-time applications.
+As a result, there are now many deep learning approaches that build on this and propose new, improved models or training procedures [6, 21, 15, 20, 32, 19, 8].
+U-Net++ was introduced in 2019 [38], and in 2023, an improved U-Net incorporating an attention mechanism was proposed [1].
+However, most of these variants require increased computational resources and are therefore not suitable for edge platforms.
+One of the most advanced real-time capable models is UNext [2], but its reliance on attention mechanisms introduces architectural complexity that currently hinders edge deployments.
+More recent work explores Mamba- and state-space-model-based crack segmentation [22], which reduces the parameter count further.
+In our work, we instead start from a standard U-Net and recover the accuracy of compact variants via knowledge distillation, an approach that has also been shown to be effective for crack segmentation in [35].
+II-B Semantic Segmentation on Edge Platforms
+In recent years, edge platforms have gained significant traction for crack segmentation.
+Some authors propose specifically designed segmentation models for edge platforms.
 For example, [3] presents MCFF-L Net - an architecture with 1.18 million parameters - combined with knowledge distillation, which achieves good segmentation results and is able to run in real-time on a Jetson Xavier NX.
 Similarly, [37] considered crack segmentation on devices for structural health monitoring.
-A related approach by [32] uses a special KD approach for measurement tasks on embedded systems.
-Other works focus on simplifying UNet for microcontroller-class hardware, like [8], which analyzed small CNN-based crack segmentation models on low-power microcontrollers.
-In contrast, [9] showed that UNet models can be compressed to meet the requirements of very resource-constrained devices such as the STM32 platform.
+A related approach by [36] uses a special KD approach for measurement tasks on edge systems.
+Other works focus on simplifying U-Net for microcontroller-class hardware, like [9], which analyzed small CNN-based crack segmentation models on low-power microcontrollers.
+In contrast, [10] showed that U-Net models can be compressed to meet the requirements of very resource-constrained devices such as the STM32 platform.
 II-C Semantic Segmentation on FPGA
@@
 Another approach to shrinking the model is about model compression and low-bit quantization.
-In [23], DeepLabV3+ is pruned using a hardware-aware genetic algorithm, while in [22] a UNet is quantized to 3-bit precision and implemented on a dedicated on-chip pipeline.
+In [26], DeepLabV3+ is pruned using a hardware-aware genetic algorithm, in [25] a U-Net is quantized down to 4-bit and implemented using a shift based accelerator, while in [24] a U-Net is quantized to 3-bit precision and implemented on a dedicated on-chip pipeline.
 Similar OpenCL- or HLS-based encoder-decoder accelerators further show the impact of quantization and optimized data flow on performance [34].
 However, in the context of crack segmentation, FPGA-based solutions are less common.
-The work in [5] represents one of the few examples, implementing a particle-filter–based crack detector directly on an FPGA for real-time inspection.
+One of the few works in this direction is [5] , which implements a particle-filter-based crack detector directly on an FPGA for real-time inspection.
 Although they use a totally different approach than we do, their approach already shows the relevance of FPGAs for building inspections.
@@
 To address this limitation, we propose offloading skip connection feature maps to external memory.
-While this introduces additional memory accesses, the resulting access pattern is largely sequential and predictable, allowing efficient utilization of memory bandwidth.
-Consequently, the associated energy overhead is expected to be limited, while significantly relaxing on-chip memory constraints and enabling the deployment of larger U-Net variants.
+This adds extra memory accesses, but the access pattern is mostly sequential and predictable, so we can use the memory bandwidth efficiently.
+We therefore expect only a small energy overhead, while we can free up enough on-chip memory to fit larger U-Net variants.
 III Methodology
-III-A Overview
-Our objective is to enable efficient crack segmentation on resource-constrained edge devices, including embedded AI accelerators and FPGAs, without compromising accuracy.
-We systematically derive a family of U-Net variants by scaling the number of feature channels, creating models spanning three orders of magnitude in parameter count.
-To mitigate the performance degradation of smaller models, we leverage knowledge distillation from a high-capacity Transformer teacher, allowing compact U-Nets to retain segmentation accuracy close to the baseline.
-This approach, combined with hardware-aware optimization, facilitates the identification of Pareto-optimal models that balance efficiency and performance across diverse edge platforms.
-III-B U-Net Channel Scaling
-We use the original UNet model [25] as a basis and generate four smaller variants by reducing the number of feature maps.
+III-A U-Net Channel Scaling
+We use the original U-Net model [28] as a basis and generate four smaller variants by reducing the number of feature maps.
 For this purpose, we introduce a parameter $c\in\{2,4,8,16,32\}$ which controls the number of all feature maps in the network while preserving the original encoder–decoder architecture.
-Setting $c=32$ reproduces the original UNet architecture.
+Setting $c=32$ reproduces the original U-Net architecture.
 This parametrization allows us to analyze how the number of feature maps influences the segmentation results as well as the runtime and power consumption.
@@
 All convolutions in the model are implemented without bias terms, as each is directly followed by a Batch Normalization layer.
-III-C Dataset and Preprocessing
-In this work, we use the CrackVision12K dataset introduced by Goo et al.
-[11].
+III-B Dataset and Preprocessing
+In this work, we use the CrackVision12K dataset [12].
 CrackVision12K is a curated benchmark for crack segmentation that combines images from several previously published crack datasets into a unified dataset with improved and standardized annotations.
@@
 In total, the dataset contains 12000 image-mask pairs.
-To ensure comparability with prior work on CrackVision12K, we use the train/validation/test split provided by Goo et al.
-[11].
+To ensure comparability with prior work on CrackVision12K, we use the train/validation/test split provided in [12].
 The training set contains 9600 images, while the validation and test sets each contain 1200 images.
-III-D Data Augmentation
+III-C Data Augmentation
 We also applied data augmentation methods during training to further improve the segmentation results and make the models more robust for practical applications.
@@
 The augmentation pipeline includes additive image noise, motion blur, horizontal flipping, and small random rotations.
-III-E Knowledge Distillation
+III-D Knowledge Distillation
 To improve the segmentation results even more, we use KD, where a stronger teacher network guides the training of a smaller student network [16].
@@
 where $\alpha\in[0,1]$ controls the trade-off between teacher supervision and supervision from the ground-truth labels.
-III-F Training Setup and Hyperparameter Search
-We first train all five U-Net variants by using supervised-learning to get the baseline models.
-In this setup, we use a combination of random search and an evolutionary algorithm as a hyperparameter search, adopted from Tschöpe et al.
-[27].
-The optimized hyperparameters include the initial learning rate of the AdamW optimizer [21], the learning rate scheduler decay factor $\gamma$, the scheduler step size, the momentum values of the AdamW optimizer, the weight decay, and the class-weights that are used in the cross-entropy loss.
-However, for testing we use fixed class-weights, computed by the pixel distribution.
-We explain this in more detail in section III-G.
+III-E Training Setup and Hyperparameter Search
+We first train all five U-Net variants with standard supervised learning to obtain the baselines.
+Then we optimize Hyperparameters with a combination of random search and an evolutionary algorithm, adopted from [30].
+The search covers the AdamW [23] learning rate, momentum parameters, and weight decay, the scheduler decay $\gamma$ and step size, and the training-time class weights.
+For KD training, we additionally optimize the temperature $T$, the mixing factor $\alpha$, and the Dice weight $\lambda_{Dice}$.
+Training-time class weights are not used at evaluation.
+Instead, we use the fixed weights described in Section III-F for all reported numbers.
 For training with KD, we additionally optimize the KD temperature $T$, the KD mixing factor $\alpha$, and the weighting factor of the Dice term $\lambda_{Dice}$.
@@
 For each model configuration, we repeat the training with five random seeds and report the mean and standard deviations of the evaluation metrics.
-III-G Evaluation Metrics
-We report both mean IoU and a weighted IoU. Let $\mathrm{IoU}_{bg}$ and $\mathrm{IoU}_{crack}$ denote the IoU values for the background and crack classes, respectively. The mean IoU is defined as
+III-F Evaluation Metrics
+Crack segmentation is in general a task with a highly imbalanced class distribution, as cracks are often very thin and small and occur much less frequently than the background.
+This is also the case with the CrackVision12K dataset that we use.
+If we only used the mean IoU as a metric, we would weight the background class and the much more interesting crack class equally, which would end in a deceptively high evaluation score.
+Therefore, in addition to mean IoU, we also use weighted IoU.
+Below, we present the definitions of both metrics in more detail.
+Let $\mathrm{IoU}_{bg}$ and $\mathrm{IoU}_{crack}$ denote the IoU values for the background and crack classes, respectively. The mean IoU is then defined as:
 $\mathrm{mIoU}=0.5\cdot\mathrm{IoU}_{bg}+0.5\cdot\mathrm{IoU}_{crack}$ (5)
-Thus, this metric weights the background pixels and crack pixels equally by a factor of $0.5$.
-Since crack pixels are much rarer but more important than background pixels in our application, we additionally report a weighted IoU:
-$\mathrm{wIoU}=w_{bg}\cdot\mathrm{IoU}_{bg}+w_{crack}\cdot\mathrm{IoU}_{crack},$ (6)
+The weighted IoU instead emphasizes the crack class more strongly than the background class:
+$\mathrm{wIoU}=w_{bg}\cdot\mathrm{IoU}_{bg}+w_{crack}\cdot\mathrm{IoU}_{crack}$ (6)
 with
 $w_{bg}=0.068,\qquad w_{crack}=0.932.$ (7)
-The weights are computed from the inverse class frequencies of the training set and normalized to sum to one.
-As a result, the weighted IoU emphasizes the crack class much more strongly than the background class.
-Therefore, a model can achieve a high mean IoU due to very good background segmentation while still obtaining a noticeably lower weighted IoU if its crack segmentation remains challenging.
+We compute the weights from the inverse class frequencies of the training set and normalize them so that they sum to one.
+The weights are fixed for all our experiments and are therefore not part of the hyperparameter search.
+A large gap between mean IoU and weighted IoU of a model means that most of its mean IoU comes from the background pixels and not from the crack pixels themselves.
 IV Implementation
-This work investigates the performance of efficient U-Net implementations on a diverse range of embedded computing platforms.
-Besides FPGA, the comparison encompasses conventional CPU, GPU, and a dedicated AI accelerator, allowing for a comprehensive evaluation of deployment strategies under different computational and energy constraints.
+In this paper we investigate the performance of U-Net implementations on four representative edge computing platforms: a conventional CPU, an edge GPU, a dedicated AI accelerator (TPU), and a custom FPGA design.
+The selected platforms cover a wide range of edge hardware under different computational and energy constraints.
 IV-A Edge Platforms
-Specifically, we target: the Raspberry Pi 5 Model B (8 GB) featuring a quad-core Cortex‑A76 CPU, the Seeed Studio reComputer J3011 equipped with an NVIDIA Jetson Orin Nano (8 GB), the Coral Edge TPU M.2 module hosted on a Raspberry Pi 5.
-For CPU-based execution, the U-Net models were run using the ONNX Runtime with floating-point weights and activations.
-On the GPU platform, two inference configurations were explored: one fully relying on floating-point precision and the other employing int8 quantization for network parameters while retaining floating-point activations.
-The GPU experiments were conducted using the TensorRT execution engine to maximize efficiency.
-On the dedicated AI accelerator, the TensorFlow Lite models, quantized to int8 for both weights and activations, were compiled using the Edge TPU compiler.
-Inference latency was measured for batch size 1 and reported as a mean value of five execution runs on a subset of 1024 input images.
-The reported latency includes data transfer times between the host and the device, as well as the time required for conversions between fp32 and int8 data formats for inputs and outputs, respectively.
-Power consumption was monitored throughout the experiments using an external USB POWER‑Z KM003C KT002 multimeter.
-Both idle and runtime conditions were evaluated, with the latter involving continuous inference execution for a duration of 20 seconds.
-The mean power values were derived from three independent measurements.
-For the Coral Edge TPU setup, power readings correspond to the complete configuration, capturing the consumption of both the M.2 module and the Raspberry Pi 5 host.
-Active cooling solutions were employed for all Raspberry Pi devices, while the Coral Edge TPU operated with an aluminum heat sink to ensure thermal stability during execution.
+We consider three edge platforms: The Raspberry Pi 5 Model B (8 GB) with a quad-core Cortex-A76 CPU, the Seeed Studio reComputer J3011 with an NVIDIA Jetson Orin Nano (8 GB), and the Coral Edge TPU M.2 module hosted on a Raspberry Pi 5.
+On the CPU we run the U-Net via ONNX Runtime in fp32.
+On the GPU we use TensorRT in two configurations: full fp32, and int8-weights with fp32-activations.
+On the Edge TPU we deploy fully int8-quantized TensorFlow Lite models compiled with the Edge TPU compiler.
+To measure inference latency, we use a batch size of 1 and calculate the average across five execution runs on a subset of 1,024 input images, including host–device transfer and fp32/int8 conversion overhead.
+Power is measured with an external USB POWER-Z KM003C KT002 multimeter, both at idle and during 20 s of continuous inference, averaged over three measurements.
+For the Coral Edge TPU, the measurement covers both the M.2 module and its Raspberry Pi 5 host.
+All Raspberry Pi devices use active cooling; the Edge TPU uses an aluminum heat sink.
 IV-B Custom Hardware Implementation on FPGA
-The FPGA implementation is built upon a High-level Synthesis (HLS) hardware library that provides custom hardware architectures for a wide range of Deep Neural Network (DNN) layers and components.
-These architectures are highly customizable, enabling the implementation of diverse neural topologies and facilitating rapid design-space exploration.
-The hardware library is implemented as a collection of C++ template functions equipped with HLS annotations.
-The overall architecture is optimized for low-power consumption and ultra-low latency operation.
-This is primarily accomplished by storing all weights and the majority of intermediate results in on-chip memory, as off-chip data transfers incur higher energy costs and introduce additional latency.
+Our FPGA implementation uses an High-level Synthesis (HLS) hardware library that provides custom hardware architectures for many common Deep Neural Network (DNN) layers and components.
+The library is built as a set of C++ template functions with HLS annotations, which makes it easy to support different network topologies and to explore the design space quickly.
+We design the architecture for low power and low latency.
+To this end, we keep all weights and most intermediate results in on-chip memory, since off-chip transfers cost more energy and add latency.
 Nevertheless, a design relying solely on on-chip resources inherently limits the size of DNN models that can be feasibly implemented.
@@
 V Results
-V-A Accuracy–Efficiency Trade-off Analysis
+V-A Scope and Fairness of the Cross-Platform Comparison
+The comparison in Table IV is a practical evaluation rather than a strictly controlled evaluation.
+The used models differ across platforms in three ways, each of which is due more to platform-specific constraints rather than to the design itself.
+First, models targeted at the Coral Edge TPU replace transposed convolutions with nearest-neighbor upsampling to avoid the accuracy drop we observed on this device; All other platforms retain the original transposed-convolution decoder.
+Second, the numerical precision is not uniform: the RPi5 runs fp32 weights and activations, the Jetson Orin Nano is evaluated in both fp32 and a mixed int8-weights/fp32-activations configuration (native to TensorRT), the Edge TPU and the FPGA executes fully int8.
+Third, the measurement instrumentation differs between platforms.
+We therefore interpret the numbers as realistic deployment results for each platform in its best-supported configuration, rather than as an apples-to-apples comparison of the same model across devices.
+V-B Accuracy-Efficiency Trade-off Analysis
 Table I summarizes the segmentation performance of all models tested in the CrackVision12K dataset before quantization.
@@
 Scaling down to $c=2$, however, incurs steep accuracy costs: a drop of 13.60pps in weighted IoU and 10.46pps in mean IoU compared to $c=16$, confirming that $c=16$ and $c=8$ represent the best efficiency–accuracy trade-offs in this family.
-As expected, SegFormer B5 outperforms all UNet variants without knowledge distillation, owing to its considerably larger capacity and modern transformer-based architecture.
-With knowledge distillation, all UNet variants improve.
+As expected, SegFormer B5 outperforms all U-Net variants without knowledge distillation, owing to its considerably larger capacity and modern transformer-based architecture.
+With knowledge distillation, all U-Net variants improve.
 The baseline model ($c=32$) achieves a weighted IoU of $51.06\%$ and a mean IoU of $71.92\%$—within 2.07pps of the teacher SegFormer B5 ($53.13\%$ weighted IoU), showing that KD brings the student nearly to the teacher’s level.
 Crucially, the efficiency gains identified without KD are preserved: $c=16$ achieves $50.63\%$ weighted IoU and $71.74\%$ mean IoU, only 0.43 and 0.18pps below $c=32$, while still benefiting from the same $4\times$ MAC reduction.
-Similarly, $c=8$ achieves $49.25\%$ weighted IoU and $70.96\%$ mean IoU—only 1.81pps and 0.96pps below the best UNet model —at a $16\times$ overall reduction in compute.
-Overall, these results show that KD not only improves segmentation results across all UNet variants, but also makes the models significantly more robust to channel reduction: the mean IoU gap between $c=32$ and $c=8$ shrinks from 2.19pps without KD to only 0.96pps with KD, which shows that KD compensates for the capacity loss introduced by pruning.
+Similarly, $c=8$ achieves $49.25\%$ weighted IoU and $70.96\%$ mean IoU—only 1.81pps and 0.96pps below the best U-Net model —at a $16\times$ overall reduction in compute.
+Overall, these results show that KD not only improves segmentation results across all U-Net variants, but also makes the models significantly more robust to channel reduction: the mean IoU gap between $c=32$ and $c=8$ shrinks from 2.19pps without KD to only 0.96pps with KD, which shows that KD compensates for the capacity loss introduced by pruning.
 In the next subsection, we evaluate each variant across several edge platforms under quantization.
 TABLE I: Comparison of U-Net variants with different channel scaling factors, with and without knowledge distillation (KD).
-The Hybrid-Segmentor model was proposed by Goo et al [11].
+The Hybrid-Segmentor model was proposed in [12].
 SegFormer B5 is included as a high-capacity reference.
@@
 Params (M)
-Hybrid-Segmentor - $63.1\pm 0.36$ - -
+Hybrid-Segmentor - $63.1\pm 0.36$ - 237.86
 SegFormer B5 $53.13\pm 0.21$ $73.42\pm 0.26$ $51.25$ $84.59$
@@
   • * This model is used as a baseline.
-V-B Hardware Implementation Results
-Table IV shows the results for knowledge-distilled models across three embedded platforms and the FPGA, reporting segmentation accuracy (weighted and mean IoU), throughput (FPS), and power and energy metrics.
+V-C Hardware Implementation Results
+As discussed in Section V-A, the configurations per platform are not identical, the comparison reflects the best-supported deployment path on each device.
+Table IV shows the results for knowledge-distilled models across three edge platforms and the FPGA, reporting segmentation accuracy (weighted and mean IoU), throughput (FPS), and power and energy metrics.
 In addition to idle and runtime power, both dynamic and runtime energy efficiency (Frames/J) are provided.
@@
 In contrast, runtime energy efficiency is affected by background processes and system-level overhead, making cross-platform comparisons less reliable.
+Runtime energy efficiency does not solve all measurement issues, but it gives a more realistic picture of how the models behave in deployment.
 Fig.
 2 further illustrates the trade-off between dynamic energy efficiency and mean IoU across the evaluated configurations.
-V-B1 FPGA results
+V-C1 FPGA results
 For an efficient FPGA implementation PTQ is used to reduce the precision of the model parameters.
@@
 In the second experiment, the weights were quantized to int4, which resulted in a considerable drop in accuracy as shown in Table II.
-We attribute this drop to PTQ and expect quantization aware training (QAT) to recover this drop.
+We see this drop as a limit of PTQ at very low bit widths and leave a follow-up with quantization aware training (QAT) to future work.
 Based on the results in Table II int8 models were chosen for implementation on FPGA, for which the AMD Vitis 2024.2 HLS tool was used and the target platform was set to TySOM-2A-ZU19EG.
 Table III presents the resources required to implement the chosen models on the target platform.
-Two experiments were conducted; in the first experiment, the resource-heavy skip connections of the UNet model were kept on chip, while in the second experiment, they were moved to off-chip memory.
+Two experiments were conducted; in the first experiment, the resource-heavy skip connections of the U-Net model were kept on chip, while in the second experiment, they were moved to off-chip memory.
 This made it possible to implement the larger model with $c=16$.
-Skip connections were moved to off-chip memory by creating an AXI interface for each skip connection, which is used for reading and writing.
-This is required due to the high read and write speeds and the fact that all skip connections are accessed in parallel.
+Skip connections were moved to off-chip memory by creating an AXI interface for each skip connection, which is used for writing and reading.
+This is required due to the high write and read speeds and the fact that all skip connections are accessed in parallel.
 On-chip buffers were added to the design to bridge the gap between the high-speed PL and the slow off-chip memory.
-The extra buffers and interfaces lead to an average increase of $2.19\%$ and $1.60\%$ in LookUp Tables (LUTs) and Flip-Flops (FFs) respectively, which is within the acceptable overhead.
+The extra buffers and interfaces lead to an average increase of 2.19pps and 1.60pps in LookUp Tables (LUTs) and Flip-Flops (FFs) respectively, which is within the acceptable overhead.
 This shows the tradeoff between moving skip connections off chip and keeping them on chip, which is less efficient for small models i.e., with $c=2$.
@@
 All FPGA implementations achieved the same IoU as the corresponding software implementations.
-V-B2 Edge Computing Platforms
+V-C2 Edge Computing Platforms
 […]
@@
 However, the model with base 32 failed to run on the Coral Edge TPU.
+$10^{0}$$10^{1}$$10^{2}$$10^{3}$696969.569.5707070.570.57171Dynamic Energy Efficiency [Frames/J]Mean IoU [%]Pareto frontBase 2Base 4Base 8Base 16Base 32RPi5Orin Nano fp32Orin Nano int8Edge TPUFPGA on-chipFPGA off-chip
+Figure 2: Pareto frontier showing the trade-off between accuracy (weighted IoU) and energy efficiency across different platforms and model sizes.
+FPGA on-chip: U-Net skip connections stored on chip, FPGA off-chip: U-Net skip connections stored off chip in the memory
 TABLE II: Post-training quantization results for models with knowledge distillation.
@@
 int8 fp32 49.06 70.88 222 6.97 12.04 43.79 18.44
-$10^{0}$$10^{1}$$10^{2}$$10^{3}$696969.569.5707070.570.57171Dynamic Energy Efficiency [frames/J]Mean IoU [%]Pareto frontBase 2Base 4Base 8Base 16Base 32RPi5Orin Nano fp32Orin Nano int8Edge TPUFPGA on-chipFPGA off-chip
-Figure 2: Pareto frontier showing the trade-off between accuracy (weighted IoU) and energy efficiency across different platforms and model sizes.
-FPGA on-chip: models UNet skip connections stored on chip, FPGA off-chip: models UNet skip connections stored off chip in the memory
 VI Conclusion and Future Work
@@
 Our results show that reducing the model size can strongly lower the computational cost, while some of the smaller variants still remain close to the baseline in terms of segmentation accuracy.
-To improve our scaled UNet models further, we applied knowledge distillation by using SegFormer B5 as the teacher model.
+To improve our scaled U-Net models further, we applied knowledge distillation by using SegFormer B5 as the teacher model.
 This improved all U-Net variants.
@@
 We also developed a custom FPGA-based architecture for single-batch inference.
-Among all tested platforms, the FPGA (on-chip) implementation with the UNet Base 4 using KD and int8 quantization achieves an energy efficiency of 204.99 Frames/J and a throughput of 398 FPS, while improving the mean IoU by 2.79pps compared to the baseline model (UNet Base 32 without KD).
-Compared to the Hybrid-Segmentor model from Goo et al.
-[11], which is - as far as we know - currently the best tested model on this dataset, our aforementioned UNet Base 4 model with KD is 6.32pps better in mean IoU.
+Among all tested platforms, the FPGA (on-chip) implementation with the U-Net Base 4 using KD and int8 quantization achieves an energy efficiency of 204.99 Frames/J and a throughput of 398 FPS, while improving the mean IoU by 2.79pps compared to the baseline model (U-Net Base 32 without KD).
+Compared to the Hybrid-Segmentor model from [12], which is - to the best of our knowledge - currently the best tested model on this dataset, our aforementioned U-Net Base 4 model with KD is 6.32pps higher in mean IoU by using the same train/val/test split.
 Therefore, our results show that the combination of model scaling, knowledge distillation, quantization, and hardware-specific implementation is useful for crack segmentation on edge devices.